SPC MODELING

Create a Control Limit chart for the following:  You are a manager at RadioTag Inc., a maker of RFID tags for US military and private use (e.g., Wal Mart).  Your process assembles radio id tags to certain specifications – 95.5 confidence intervals, however as they come off the assembly line a test is run on each tag to determine whether the frequency works or does not (hint the data are dichotomous distributed binary).  Below are the data you collected over a 2-week period:                                                                                                                                                                                                                                                                                                                                    GRAPH GOES UNDER THIS

Day sample size         non-working tags

1                    100                       15

2                    100                       13

3                    100                       12

4                    100                       11

5                    100                       16

6                    100                       19

7                    100                       5

8                    100                       8

9                    100                       19

10                  100                       15

11                  100                       2

12                  100                       6

13                  100                       9   

14 100                       10

P bar =

SD of P bar =

UCLP Bar =

LCLP bar =

B) Graph and Plot the data.  Round answers to three places.

Clearly identify the P bar, s.d. of Pbar, UCL, LCL (make sure the graph elements are labeled) and identified in A above. Place a box around these final answers.  Failure to place a box will result in failure to earn points.

C) Compare p value to a.  What is the result?
Answer #1
Day n np p = np / n
1 100 15 0.1500
2 100 13 0.1300
3 100 12 0.1200
4 100 11 0.1100
5 100 16 0.1600
6 100 19 0.1900
7 100 5 0.0500
8 100 8 0.0800
9 100 19 0.1900
10 100 15 0.1500
11 100 2 0.0200
12 100 6 0.0600
13 100 9 0.0900
14 100 10 0.1000
Average 0.1143
p_bar

p_bar = 0.1143
SD of p_bar, Sp = [p_bar * (1 - p_bar) / n]1/2 = sqrt(0.1143*(1 - 0.1143) / 100) = 0.0318

Confidence level = 95.5%, so, α = 0.045 and α/2 = 0.0225
So, zα/2 = Normsinv(1 - 0.0225) = 2.00

UCLp = p_bar + zα/2 * Sp = 0.1143 + 2*0.0318 = 0.1779
LCLp = Max(0, p_bar - zα/2 * Sp)= Max(0, 0.1143 - 2*0.0318) = 0.0507